OpenStudy (anonymous):

logbase3(1/9)=x find x please explain

OpenStudy (anonymous):

can you use equation it is hard to understand

OpenStudy (anonymous):

but the answer can be -2

OpenStudy (anonymous):

\[\log_{3}(1/9)=x \]

OpenStudy (anonymous):

No answer is -2

OpenStudy (hoblos):

x=logbase3(1/9) = logbase3(1) - logbase3(9) = 0 - logbase3(3^2) = - 2logbase3(3) = -2

OpenStudy (turingtest):

\[3^1=3\]\[3^0=1\]\[3^{-1}=\frac13\]\[3^{-2}=\frac19\]in general\[a^{-n}=\frac1{a^n}\]

OpenStudy (anonymous):

then \[\log_{3} (3^{-2}) \] \[-2\log_{3} (3)\] -2

OpenStudy (turingtest):

\[\log_3(\frac19)\]is like asking "what power do you raise 3 to to get 1/9?" the answer is -2